OpenStudy (anonymous):

A motorcycle stunts man belonging to a fair, rides over the vertical walls of a circular well at an average speed of 54 kph for 5 minutes. If the radius of the well is 5 meters then the distance traveled is

OpenStudy (anonymous):

@ParthKohli please help

OpenStudy (anonymous):

@lgbasallote please help

OpenStudy (anonymous):

will it depend on the radius of the well?

OpenStudy (lgbasallote):

sounds like circular motion....

Parth (parthkohli):

54 kph = 54000meters per hour 54000/5 = 10,800 meters.

Parth (parthkohli):

10.8 kilometers.

OpenStudy (lgbasallote):

lol that's a big well :p

OpenStudy (anonymous):

4.5 km

OpenStudy (anonymous):

no wrong ans. is 4.5 kms.

Parth (parthkohli):

Oh lol

OpenStudy (lgbasallote):

get the circumference first 54 x 1/12 hours = 4.5

OpenStudy (anonymous):

just find out the distance covered in 5 minutes. 54 km in 60 minutes so, in 1 minute, 54/60 km in 5 minutes, 54/60 * 5 km=4.5 km

OpenStudy (anonymous):

I meant ur ans is wrong ... correct ans is 4.5 kms

OpenStudy (lgbasallote):

but does the raidus even matter =))

Parth (parthkohli):

Maybe the mods will punish me.. I should answer questions I know

OpenStudy (anonymous):

i guess not, @lgbasallote :)

OpenStudy (lgbasallote):

i answer a lot of questions blindly parth :P just dont take the populated ones....

OpenStudy (lgbasallote):

too many witnesesses @ParthKohli =_=

OpenStudy (anonymous):

radius is given just for fun, i guess.. :/

OpenStudy (anonymous):

thanks @Arnab09

OpenStudy (anonymous):

you guys are fun

OpenStudy (dumbcow):

well you could add a part b) to the question: how many times did he ride around the well? divide distance by circumference: 4500/10pi = 1413.7

OpenStudy (lgbasallote):

he went around a thousand times?? ill be dizzy if i were him

OpenStudy (anonymous):

HEHEHE

OpenStudy (anonymous):

:D

OpenStudy (lgbasallote):

lol it should be 143 @dumbcow :P

Parth (parthkohli):

My profession, certainly not in circus.

OpenStudy (anonymous):

Oh this is great really very great

OpenStudy (anonymous):

Then what do u think ... people who are replying work in circus.....@ParthKohli

OpenStudy (dumbcow):

@lgbasallote , you sure? 4.5km = 4,500m

OpenStudy (anonymous):

Thanks Guys

OpenStudy (lgbasallote):

4500/10 = 450...450/3.14 is 143.3

OpenStudy (dumbcow):

haha i Fail....forgot parenthesis on 10pi in calculator

OpenStudy (lgbasallote):

:PPP

sam (.sam.):

I'm thinking of using \[v=54kph\] \[v=15ms^{-1}\] maybe you could find the angular velocity \[v=r \omega\] \[\omega=3rad/s\] -------------------------------- Then use \[\theta=\omega t+\frac{1}{2}\alpha t^2\]

sam (.sam.):

Since there's no angular acc, θ=ωt θ=(3)(300)=900radians

sam (.sam.):

Which is 900m